1. Question: Selling and administrative costs are often called __________ costs and are __________ as incurred.

Fill in the blank choices for question

Administrative

Beginning

Contribution

Cost driver

Direct labor

Direct labor-hours

Direct materials

Dollars of direct labor cost

Ending

Expensed

Fixed

Fixed manufacturing overhead

For the period ended

Income statement

Indirect labor

Indirect materials

Job cost

Level of activity

Machine-hours

Manufacturing

Manufacturing overhead

Margin

Name of company

Overapplied

Period

Predetermined overhead rate

Process cost

Selling

Statement of cost of goods manufactured

Total budgeted overhead

Underapplied

Variable costing

Work inprocess inventory

2. Question: If Cost of Goods Manufactured is $306,790, beginning Work in Process Inventory, $25,000, and Cost to Manufacture, $300,000, the ending Work in Process Inventory balance is $__________.

Fill in the blank answer choices:

$7500

$14000

$18210

$70000

$120000

$180000

$225000

$232500

$420000

3. Question: Using the same data as in the preceding question, with Direct Labor costing $140,000 and Direct Materials costing $90,000, Manufacturing Overhead is $__________.

Fill in the blank answer choices:

$7500

$14000

$18210

$70000

$120000

$180000

$225000

$232500

$420000
Answer #1

1. Selling and administrative costs are often called period costs and are manufacturing overhead as incurred.

2. 18210

25000-(306790-300000)

3. 70000

MO = Cost to manufacture - prime cost

= $300,000 - ($140,000 + $90,000)

= $70,000

Thus, manufacturing overhead is $70,000.
Answer #2


ANSWER :


1. period ;  manufacturing overhead .


2. 


Toal input to the production = Total output from the production


=> Work-in-process inventory + cost-to-manufacture = Cost of goods sold + ending work-in-process

=> 25000 + 300000 = 306790 + ending work-in-process

=> Ending work-in-process = 25000 + 300000 - 306790  = 18210 ($) (ANSWER).


3.


Cost-to-manufacture = Labor cost + Material cost + Overhead cost 

=> 300000 = 140000 + 90000 + Overhead cost

=> Overhead cost = 300000 - 140000 - 90000 = 70000 ($) (ANSWER).


answered by: Tulsiram Garg